Olivia has $30 more than Carl. if they each had $7 less, the sum of the funds would equal the amount that olivia has now. how much money does each have now........ please show steps!!
Found 3 solutions by Boreal, fractalier, solver91311:
Answer by Boreal(15235)

Carl has x
Olivia has x+30
If Olivia had x+30-7, or x+23, Carl would have x-7
Add those, x+23 and x-7, and they equal what Olivia has now, or x+30
2x+16=x+30
subtract x from both sides and subtract 16 from both sides
x=$14
Carl has $14, and Olivia has $44.
If they each had $7 less, Carl would have $7 and Olivia $37. That adds to $44, which is what Olivia has.

Answer by fractalier(6550)

Call what Olivia and Carl have, x and y. Thus we have
x = y + 30 and
(x - 7) + (y - 7) = x
A few ways to solve this, but let us substitute the first equation into the second and we get
(y+30 - 7) + (y - 7) = y + 30
2y + 16 = y + 30
Now subtract y from both sides, then subtract 16...
y + 16 = 30
y = $14 for Carl
x = y + 30 = $44 for Olivia

Answer by solver91311(24713)



Let represent Olivia's amount and represent Carl's amount.

We are given that:



and



Solve the 2X2 system for and

John

My calculator said it, I believe it, that settles it